Calulatete the net calorific value of a coal sample having the following composition: C-80%, H=5%, O = 4%, N= 3%, S = 3.5% and ash= 5%.


1
Expert's answer
2021-06-18T07:32:33-0400

C-80%, H=5%, O = 4%, N= 3%, S = 3.5% and ash= 5%.


Solution:

Dulong’s formula can be used:

HCV = (1/100)[8080C + 34500(H-O/8) + 2240S],

where C, S, O, H are the percentages of carbon, sulphur, oxygen and hydrogen.

So, substitute all the given values in the formula and calculate so that you will get HCV:

HCV = (1/100) × [8080×80 + 34500×(5 - 4/8) + 2240×3.5] = 8094.9 cal/g


Apply the formula for net calorific value (NCV): NCV = (HCV-0.09H*587)

Therefore,

NCV = (8094.9 - 0.09×5×587) = 7830.75 cal/g


Answer: Net calorific value (NCV) of a coal is 7830.75 cal/g.